OpenStudy (nightmarenight):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 98

OpenStudy (anonymous):

49

OpenStudy (anonymous):

\[ x^2 =98\\ x=\pm 7\sqrt 2 \]

OpenStudy (anonymous):

\[ x=\pm9.9 \]

OpenStudy (texaschic101):

x^2 = 98 x = sq rt 98 x = +- 9.89 rounds to +- 9.9

OpenStudy (nightmarenight):

Thank you @eliassaab

OpenStudy (anonymous):

YW